If you have a vertical string instead of a horizontal would the object travel the same distance w/ the same speed? What vectors of forces would be involved?
Answer #1

Solution:

If there is a vertical string , the forces acting on it will be gravity , and the tension in the string . This results in a greater force when it accelerates downwards because there will be the force due to gravity along with the accelerating force.

For an object hanging from a vertical string,

Tension force = T -mg = ma

T-mg ma ng when the acceleration is downwards T-mg=ma and the tension in the string will be T-mg =-ma